Determine two coterminal angles (one positive and one negative) for the given angle. Give your answers in radians.

a) 8pie/9 b)8pie/45

I know that you have to add 2pie but I do not know how to do that.
Answer #1
The last two should read:
π/2 + 2π = (5/2)π
π/2 + 4π = (9/2)π
answered by: Lynne
Answer #2
Coterminal angles for an angle α are angles of the form α±2kπ where k is different from zero.
For example, some coterminal angles of π/2 are:
π/2 - 4π = -(7/2)π
π/2 - 2π = -(3/2)π
π/2 + 2π = (3/2)π
π/2 + 4π = (7/2)π
answered by: Ch